OpenStudy (anonymous):

PLEASE HELPPPPP MEEEEEEEEE !!! Cathy's salary is 3/7 of Nora's salary and is 5/4 of Teresa's salary. Nora's salary is what fraction of Teresa's salary?

OpenStudy (anonymous):

can you help explain why i wanna know HOT to do the problem more than just gettin an answer =/

OpenStudy (anonymous):

HOW*

OpenStudy (anonymous):

It's like this: "X" has 1/2 of "N" 's salary. What would X's salary be if N was 10 ?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

Alright, now what if we didn't know N's salary. X has 1/2 of N, N has 1/2 of W's salary. W's salary is 40 What's N's salary? Then find X's after you find N's.

OpenStudy (anonymous):

20 and X is 10

OpenStudy (anonymous):

Alright. So you understand how to get that. It's essentially the same concept. You got 20 and 10 by first multiplying W by 1/2 , then you multiplied 20 by 1/2 and got 10 right? Now what if you wanted to skip N and go straight to X's value?

OpenStudy (anonymous):

kinda lost now tbh

OpenStudy (anonymous):

It would be the same order, X = W * 1/2 * 1/2 So, for your problem it would be: Cathy = Teresa * 5/4 * 3/7

OpenStudy (anonymous):

am I dividing?

OpenStudy (anonymous):

wait why would you multiply?

OpenStudy (anonymous):

It's multiplication, but with fractions. W = 40 N = W * 1/2 X = N * 1/2 And you get 10

OpenStudy (anonymous):

i seriously dont get why i have to multiply.....

OpenStudy (anonymous):

What's 1/2 * 1/2 ?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

.0025

OpenStudy (anonymous):

Not exactly..

OpenStudy (anonymous):

.25

OpenStudy (anonymous):

Yes! So 1/2 * 1/2 = .25 OR 1/4 And what does 40 * 1/4 equal?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

Yes! And that's why you multiply to get your answer. We multiplied the fractions 1/2 * 1/2 because N = 1/2 * W X = 1/2 * N If you think about it, we MULTIPLIED W TWICE by 1/2. It looks confusing because 1/2 * W appears once... But if you think about it N = 1/2 * W and X = 1/2 * N ... But actually if you replace N with( 1/2 * W ) what do you get?

OpenStudy (anonymous):

Replace N in X to what it's equal too.

OpenStudy (anonymous):

... wow this is seriously confusing

OpenStudy (anonymous):

Sorry, I'll try a different approach.

OpenStudy (anonymous):

Cathy's salary is 3/7 of Nora's salary Cathy = 3/7 * Nora Nora is 5/4 of Teresa's salary Nora = 5/4 * Teresa Now what would Cathy's salary be if Teresa's salary was 20$?

OpenStudy (anonymous):

Refer to the Mathematica attachment.

OpenStudy (anonymous):

k thats even more confusing

OpenStudy (anonymous):

why is it 7/3 x 5/4

OpenStudy (anonymous):

thats what the answer says it is

OpenStudy (anonymous):

cuz im dividing no?

OpenStudy (anonymous):

Let c, n and t be Cathy's, Nora's and Teresa's salary respectively. Cathy's salary is 3/7 of Nora's salary. c = 3/7 n In a similar fashion c = 5/4 t If c = a and c = b then "a" must equal "b" 3/7 n = 5/4 t Can you go ahead and solve the above for "n" ?

OpenStudy (anonymous):

Multiplying fractions is like dividing ... Except you multiply by the numerator and divide by the denominator. But when you have to fractions it's easier to find a common denominator by cross multiplication. Then with *that* fraction, you can say that's Noras salary as a fraction of Teresa's

OpenStudy (anonymous):

I guess @yomamabf bailed out.

OpenStudy (anonymous):

35/12 x t

OpenStudy (anonymous):

no im still here im just tryna figure this out sorry im really bad at math

OpenStudy (anonymous):

Example 1 \[1/2 * 10 = 10/2\] Example two: \[3/5 * 10 = 30/5 \] See, there is some division, but you have to multiply by the top and divide by the bottom.

OpenStudy (anonymous):

What you are doing though is multiplying fractions. https://www.khanacademy.org/math/pre-algebra/fractions-pre-alg/multiplying-fractions-pre-alg/v/multiplying-fractions

OpenStudy (anonymous):

wait cant it be 5/4n = 3/7

OpenStudy (anonymous):

3/7 n = 5/4 t Divide each side by 3/7 (3/7)/(3/7) n = (5/4)/(3/7) t n = (5/4)/(3/7) t n = (5/4)*(7/3) t n = (5/4)*(7/3) t n = 35/12 t

OpenStudy (anonymous):

Nora = 5/4 * Teresa Cathy = 3/7 * 5/4 * Teresa You don't know what Teresa is. Combine what you can -- This is also called "Like Terms" Cathy = 3/ 7 * 5 / 4 - You got rid of Teresa because it just wants to know the fractional part. Just take the khan academy link on multiplying fractions and you'll have our answer.

OpenStudy (anonymous):

okay im spending way too much time on this okay thanks guys!

OpenStudy (anonymous):

You're welcome. Sorry I couldn't explain better. Hopefully in the future I can give more thorough explanations.

OpenStudy (anonymous):

thanks anyways you're awesome i'm gonna ask another question